3. Donut Dolly purchased a single share of Krispy Kreme Donuts six years ago for $7.78. In addition to the purchase price, a $1.50 transaction fee was charged at the time of purchase. Today the stock is selling for $19.31. If Dolly sells the stock today and pays no sell fee, what rate of return will she earn? The stock paid no dividend while Dolly owned it. [Hint: Any type of fee you pay during purchase, gets added to original purchase value. Fee during sale gets subtracted from sale value. We use the net values (added and/or, subtracted).]
Answer #1

Solution: -

Present value = Purchase price of donuts six years ago + transaction fee = 7.78 + 1.50 = 9.28

Future value = 19.31

n = 6 years

r =?

FV = PV * (1 + r) ^ n

19.31 = 9.28 * (1 + r) ^ 6

19.31 / 9.28 = (1 + r) ^ 6

2.080 = (1 + r) ^ 6

(2.080) ^ 1 / 6 = 1 + r

1.129898 = 1 + r

r = 1.129898 - 1

r = 0.129898 or 12.98% or 13% (rounded off)